# Misaki Kojima    Naoki Nishida

### v1

@@ 6 Reduction of Liveness Properties to APR Problems
Informally speaking, liveness properties state that “something good” will happen in the future [BK08, Section 3.4]. As a liveness property w.r.t. a system, a source set $P$, and a set $Q$ for “something good”, we consider the property that any (possibly non-terminating) execution of the system starting from any state in $P$ reaches a state in $Q$. This kind of safety properties is formulated for ARSs as follows:
@@
Unfortunately, this problem cannot be reduced to partial validity of $P\Rightarrow Q$ w.r.t. $\mathcal{A}$, while $P\Rightarrow Q$ seems natural for liveness properties. The shortcoming is that partial validity does not take into account any infinite execution path. In this section, we introduce a stronger validity, called total validity, which takes into account all (possibly infinite) execution paths. Then, for a partially valid APR predicate with a cyclic-proof tree, we show a necessary and sufficient condition for the tree to ensure total validity, showing how to apply APR analysis to verification of liveness properties.
@@ 6.2 A Criterion for Total Validity of Partial Valid APR Predicates
Let $((V,\mbox{\sl a},\mbox{\sl r},\mbox{\sl p}),\xi)$ be a partially valid APR proof of an APR predicate $P\Rightarrow Q$ w.r.t. $\mathcal{A}$, and $(V^{\prime},E)$ the proof graph of $((V,\mbox{\sl a},\mbox{\sl r},\mbox{\sl p}),\xi)$. Then, $(V^{\prime},E)$ is acyclic if and only if $\mathcal{A}\mathrel{\mathrel{\models^{\forall}_{\mathrm{total}}}}P\Rightarrow Q$.
Proof 6.4 (Proof (Sketch)).
Both the only-if and if parts are proved by contradiction. For the if part, we consider a bud node $v$ and its companion $v^{\prime}$ in a cycle such that $\xi(v)=v^{\prime}$. By the assumption, we have that $\mbox{\sl a}_{L}(v)=\mbox{\sl a}_{L}(v^{\prime})$. It follows from Proposition˜4.6 23 that for any element $s\in\mbox{\sl a}_{L}(v)$, there exists an element $s^{\prime}\in\mbox{\sl a}_{L}(v^{\prime})=\mbox{\sl a}_{L}(v)$ such that $s^{\prime}\to_{\mathcal{A}}^{+}s$ and the reduction has no element in $Q$. Using this claim, we can show an infinite execution path that starts with an element in $P$ and does not include any element in $Q$. This contradicts the assumption that $\mathcal{A}\mathrel{\mathrel{\models^{\forall}_{\mathrm{total}}}}P\Rightarrow Q$.
We call a closed APR pre-proof acyclic if the proof graph of the APR pre-proof is acyclic. As a consequence of Theorem˜6.3, for an APR predicate $P\Rightarrow Q$, the existence of an acyclic APR pre-proof is a sufficient condition for total validity of $P\Rightarrow Q$ w.r.t. $\mathcal{A}$.
@@
Once we construct an APR proof, the total validity is decidable.
Corollary 6.6.
Let $((V,\mbox{\sl a},\mbox{\sl r},\mbox{\sl p}),\xi)$ be an APR proof of a partially valid APR predicate $P\Rightarrow Q$ w.r.t. $\mathcal{A}$. Then, both of the following statements hold:
  1. 1.
    $\mathcal{A}\mathrel{\mathrel{\models^{\forall}_{\mathrm{total}}}}P\Rightarrow Q$ if and only if the proof graph of $((V,\mbox{\sl a},\mbox{\sl r},\mbox{\sl p}),\xi)$ is acyclic, and
  2. 2.
    total validity of $P\Rightarrow Q$ w.r.t. $\mathcal{A}$ is decidable.
Proof 6.7.
The first claim follows Theorem˜6.3. It is decidable whether a finite graph is acyclic. Thus, the second claim holds.
@@
Let us continue with Example˜6.2. We have an acyclic APR proof for the APR predicate (12). Therefore, by Corollary˜6.5, the APR predicate (12) is totally valid w.r.t. $\mathcal{A}_{2}$, and thus the AITS is starvation free for $P_{0}$.
    (Axiom) (17)  (Subs) (13)        (Axiom) (17)  (Subs) (18)  (Der) (14)†        (Axiom) (17)  (Subs) (21)   (14)†  (Der)     (19)      (Der)     (15)        (Axiom) (17)  (Subs) (20)   (14)†  (Der)     (16)  (Der)                  (12)
@@ 7 Related Work
The work in this paper is the adaptation of the APR frameworks in [CL18, KN23jlamp] to ARSs, providing an abstract foundation for APR analysis. $\mathbb{K}$ framework [RS10] is a more general setting of rewriting than LCTRSs, and the proof system for all-path reachability has been implemented in $\mathbb{K}$ [SCMMSR14, SCMMSR19]. The race-freedom of Peterson’s mutual exclusion algorithm has been proved in [SCMMSR19] by means of the APR approach, while starvation freedom is not considered. A comparison of all-path reachability logic with CTL* can be seen in [SCMMSR19]. Our previous work [KN23padl, KN24rp] for the verification of starvation freedom by means of APR analysis with LCTRSs reduces starvation freedom to safety APR predicates, and does not use the approach to liveness properties in this paper.
@@
The cyclic-proof system for partial validity of APR predicates in this paper is a simplified adaptation of the well-known cyclic-proof sytem $\mbox{CLKID}^{\omega}$ [Bro06phd, Chapter 5]. Compared with sequents, APR predicates are very simple, and our APR proofs always satisfy the global trace condition.
@@ 8 Conclusion
In this paper, we first reformulated inference rules for partial validity of APR predicates w.r.t. ARSs, and adapted the cyclic-proof system for partial validity w.r.t. LCTRSs to ARSs. Then, we showed how to apply APR analysis to safety verification. Finally, we introduced total validity of APR predicates w.r.t. ARSs and we showed if there is an acyclic APR proof for an ARP predicate w.r.t. an ARS, then the APR predicate is totally valid w.r.t. the ARS.
Our APR pre-proofs do not consider the generalization for bud nodes. To be more precise, in the current definition, a bud and its companion have the same APR predicate. As described in Footnote˜4, the relaxation of bud-companion relationships—an APR predicate can be more general than that of its companion—is one of our future work.
@@ Appendix A Omitted Proofs
We first show a naive construction of APR pre-proofs of $P\Rightarrow Q$, where we do not split any source sets. It follows from Proposition˜3.9 that for each APR predicate, there is exactly one applicable rule in $\mathsf{DVP}_{\bot}(\mathcal{A})$. Let $v_{0}$ be the root of a derivation tree we construct now, where $\mbox{\sl a}(v_{0})=(P\Rightarrow Q)$. In the following, we use $V$ as a set of nodes that have already been applied a rule in $\mathsf{DVP}_{\bot}(\mathcal{A})$ or considered a bud node; we use $V_{\mathrm{Der}}$ as a set of nodes that can be companions of other nodes; we use $U$ as a queue of nodes that have not been applied any rule in $\mathsf{DVP}_{\bot}(\mathcal{A})$ yet. Then, we construct an APR pre-proof by breadth-first search (cf. [KN24jip]):

### v2

@@ 6 Reduction of Liveness Properties to APR Problems
Informally speaking, liveness properties state that “something good” will happen in the future [BK08, Section 3.4]. As a liveness property w.r.t. a system, a source set $P$, and a set $Q$ for “something good”, we consider the property that any (possibly non-terminating) execution of the system starting from any state in $P$ reaches a state in $Q$. This kind of safety properties is formulated as a problem for ARSs as follows:
@@
Unfortunately, this problem cannot be reduced to partial validity of $P\Rightarrow Q$ w.r.t. $\mathcal{A}$, while $P\Rightarrow Q$ seems natural for liveness properties. The shortcoming is that partial validity does not take into account any infinite execution path. In this section, we introduce a stronger validity, called total validity, which takes into account all (possibly infinite) execution paths. Then, for a partially valid APR predicate with an APR proof, we show a necessary and sufficient condition for the tree to ensure total validity, showing how to apply APR analysis to verification of liveness properties.
@@ 6.2 A Criterion for Total Validity of Partial Valid APR Predicates
Let $((V,\mbox{\sl a},\mbox{\sl r},\mbox{\sl p}),\xi)$ be a partially valid APR proof for an APR predicate $P\Rightarrow Q$ w.r.t. $\mathcal{A}$, and $(V^{\prime},E)$ the proof graph of $((V,\mbox{\sl a},\mbox{\sl r},\mbox{\sl p}),\xi)$. Then, $(V^{\prime},E)$ is acyclic if and only if $\mathcal{A}\mathrel{\mathrel{\models^{\forall}_{\mathrm{total}}}}P\Rightarrow Q$.
Once we construct an APR proof, the total validity is decidable.
Corollary 6.4.
Let $((V,\mbox{\sl a},\mbox{\sl r},\mbox{\sl p}),\xi)$ be an APR proof for a partially valid APR predicate $P\Rightarrow Q$ w.r.t. $\mathcal{A}$. Then, total validity of $P\Rightarrow Q$ w.r.t. $\mathcal{A}$ is decidable.
Proof 6.5.
It is decidable whether a finite graph is acyclic, and thus, by Theorem˜6.3, totally validity is decidable.
We call a closed APR pre-proof acyclic if its proof graph is acyclic. As a consequence of Theorem˜6.3, for an APR predicate $P\Rightarrow Q$, the existence of an acyclic APR pre-proof is a sufficient condition for total validity of $P\Rightarrow Q$ w.r.t. $\mathcal{A}$.
@@
Let us continue with Example˜6.2. We have an acyclic APR proof for the APR predicate (12) shown in Figure˜8. Therefore, by Corollary˜6.6, the APR predicate (12) is totally valid w.r.t. $\mathcal{A}_{2}$, and thus the AITS is starvation free for $P_{0}$.
         (Axiom) (17)  (Subs) (13)        (Axiom) (17)  (Subs) (18)  (Der) (14)†        (Axiom) (17)  (Subs) (21)   (14)†  (Der)     (19)      (Der)     (15)        (Axiom) (17)  (Subs) (20)   (14)†  (Der)     (16)  (Der) (12) $\{\langle\,\mathsf{wait}_{0},p_{1},\sigma\,\rangle\mid p_{1}\in\mathit{State}_{1},\sigma\in\mathcal{S},\sigma(b_{0})\,{=}\,\mathsf{true}\}\Rightarrow\{\langle\,\mathsf{crit}_{0},p_{1},\sigma\,\rangle\mid p_{1}\in\mathit{State}_{1},\sigma\in\mathcal{S}\}$
@@ 7 Related Work
The work in this paper is the adaptation of the APR frameworks in [CL18, KN23jlamp] to ARSs, providing an abstract foundation for APR analysis. The cyclic-proof system for partial validity of APR predicates in this paper is a simplified adaptation of the well-known cyclic-proof system $\mbox{CLKID}^{\omega}$ [Bro06phd, Chapter 5]. Compared with sequents, APR predicates are very simple, and our APR proofs always satisfy the global trace condition.
Runtime-error verification by means of APR analysis w.r.t. constrained rewrite systems such as LCTRSs has been investigated. $\mathbb{K}$ framework [RS10] is a more general setting of constrained rewriting than LCTRSs, and the proof system for partial validity of APR predicates consisting of constrained terms has been implemented in $\mathbb{K}$ [SCMMSR14, SCMMSR19]. The race-freedom of Peterson’s mutual exclusion algorithm has been proved in [SCMMSR19] by means of the APR approach, while starvation freedom is not considered. A comparison of all-path reachability logic with CTL* can be seen in [SCMMSR19].
@@
Our previous work [KN23padl, KN24rp] for the verification of starvation freedom by means of APR analysis with LCTRSs reduces starvation freedom to safety APR predicates, and does not use the approach to liveness properties in this paper. On the other hand, the previous work [KN24rp] deals with process-fairness for starvation freedom.
@@ 8 Conclusion
In this paper, we first reformulated inference rules for partial validity of APR predicates w.r.t. ARSs, and adapted the cyclic-proof system for partial validity w.r.t. LCTRSs to ARSs. Then, we showed how to apply APR analysis to safety verification. Finally, we introduced total validity of APR predicates w.r.t. ARSs and showed if there is an acyclic APR proof for an APR predicate w.r.t. an ARS, then the APR predicate is totally valid w.r.t. the ARS.
Our APR pre-proofs do not consider the generalization for bud nodes. To be more precise, in the current definition, a bud node and its companion have the same APR predicate. As described in Footnote˜3, the relaxation of bud-companion relationships—an APR predicate can be more general than that of its companion—is one of our future work.
@@ Appendix A Omitted Proofs
We first show a naive construction of APR pre-proofs of $P\Rightarrow Q$, where we do not split any source sets. It follows from Proposition˜3.7 that for each APR predicate, there is exactly one applicable rule in $\mathsf{DVP}_{\bot}(\mathcal{A})$. Let $v_{0}$ be the root of a derivation tree we construct now, where $\mbox{\sl a}(v_{0})=(P\Rightarrow Q)$. In the following, we use $V$ as a set of nodes that have already been applied a rule in $\mathsf{DVP}_{\bot}(\mathcal{A})$ or considered a bud node; we use $V_{\mathrm{Der}}$ as a set of nodes that can be companions of other nodes; we use $U$ as a queue of nodes that have not been applied any rule in $\mathsf{DVP}_{\bot}(\mathcal{A})$ yet. Then, we construct an APR pre-proof by breadth-first search (cf. [KN24jip]):